OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. m = 3, point (5, 10)

OpenStudy (anonymous):

\[y - y1 = m(x-x1)\]

OpenStudy (ivettef365):

you use formula y = mx + b you have m = 3, and you have points (5,10) represents the x and the y, needs to find b